OpenStudy (darkprince14):

What does it mean if the product of the solutions of a sequence converges to 0?

OpenStudy (agent0smith):

Wouldn't that mean the terms of the sequence also eventually converge to zero? Since the limit of a product is the product of the limits.

OpenStudy (darkprince14):

I took the limit of the sequence, and it has a nonzero value

OpenStudy (darkprince14):

The product will be zero given certain conditions for the initial value of the sequence. Does that mean that, given these certain conditions, the limit of the sequence will also be 0?

OpenStudy (agent0smith):

What is the sequence? It sounds like you have a lot more information than we do

OpenStudy (darkprince14):

\[x_n = \frac {q}{p + x_n ^\nu}\] this equation where p, q and v are natural numbers. The condition was p > q-1

OpenStudy (agent0smith):

So people can read it \[\Large x_n = \frac {q}{p + x_n ^\nu}\] This doesn't make a lot of sense to me. It shows xn as a function of itself.

OpenStudy (darkprince14):

it was equal to x_(n+1). sorry. my bad.

OpenStudy (darkprince14):

\[\Large{x_{n+1} = \frac{q}{p + x_n^\nu}}\]

OpenStudy (agent0smith):

It seems like with the above conditions, the terms should always eventually converge to zero.

OpenStudy (darkprince14):

Oh so that's is what it is for. thanks :)

OpenStudy (agent0smith):

Actually I guess it does depend on what you choose the first x as... when I pick something like x=0.1, it doesn't approach zero